The length of a rectangle is 3 times as much as the width. the width is given as 6. find the length and the perimeter of the rectangle.
Answer by tutorcecilia(2152)

Formula for perimeter:
P=2(length + width)
P=?
Lenght = (3)(6)
Width = 6
.
Plug-in the values and solve for the perimeter:
P=2(length + width)
P=2((3)(6)+6)
P=2(18+6)
P=2(24)
P=48